# Interpretable and flexible non-intrusive reduced-order models using reproducing kernel Hilbert spaces

### v1

@@ 1 Introduction
Our proposed kernel-based non-intrusive ROMs, which we call “Kernel ROMs”, share similarities with the aforementioned approaches while overcoming some noticeable drawbacks. Like other approaches, we begin by applying POD or QM dimension reduction to a set of training snapshots and learn a function that approximates the system dynamics within a latent space. However, instead of modeling the ROM dynamics as a polynomial and learning the polynomial coefficients through least-squares regression as in OpInf and LaSDI, we use regularized kernel interpolation [36, 49, 50] to model the reduced dynamics with a function belonging to a user-defined reproducing kernel Hilbert space (RKHS). The structure of the learned function depends on the positive-definite kernel that defines the RKHS. For example, if the governing FOM has a polynomial structure, we can use a kernel induced by a feature map to compute ROM dynamics that share the same polynomial structure. On the other hand, if the FOM dynamics have unknown or only partially known structure, a more generic nonlinear kernel can be used to model the unknown part of the ROM dynamics. In this sense, our proposed approach has a natural way of incorporating closure terms into the ROM dynamics. While kernel methods have been used to emulate reduced dynamics in previous work [41, 62], these approaches are intrusive in that they assume that the FOM dynamics can be sampled explicitly, and they do not demonstrate a way to inject explicit structure into the learned ROM. The authors in [3] use kernel methods to augment a DMD model, resulting in a fully data-driven surrogate model, and implicitly inject structure by modeling nonlinear terms using polynomial kernels. While this approach is similar to ours, we focus on constructing non-intrusive ROMs, and can model nonlinear terms explicitly using feature map kernels. In summary, the proposed approach is entirely data-driven, can produce interpretable and flexible ROMs, and yields computable a posteriori error bounds between the non-intrusive ROM and FOM solutions.
@@ 2.1 Scalar-valued kernel interpolation
We begin with a review of regularized kernel interpolation for scalar-valued functions. By the Moore–Aronszajn Theorem (see, e.g., [49, Theorem 3.10]), a positive-definite kernel function defines a unique Hilbert space with desirable properties. This result leads to the Representer Theorem — the key result used for computing an optimal interpolant in an RKHS — as well as a pointwise error bound on the interpolant.
@@
A function $K:{}^{n_{x}}\times{}^{n_{x}}\to\real$ is a (real-valued) kernel function if it is symmetric, i.e., $K(\mathbf{x},\mathbf{x}^{\prime})=K(\mathbf{x}^{\prime},\mathbf{x})$ for all $\mathbf{x},\mathbf{x}^{\prime}\in{}^{n_{x}}$. A kernel function $K$ is said to be positive definite if for any matrix $\mathbf{X}=[~{}\mathbf{x}_{1}~{}~{}\cdots~{}~{}\mathbf{x}_{m}~{}]\in{}^{n_{x}\times m}$ with pairwise distinct columns, the kernel matrix $K(\mathbf{X},\mathbf{X})\in{}^{m\times m}$ with entries $K(\mathbf{X},\mathbf{X})_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$ is positive semi-definite.
@@
Let $K:{}^{n_{x}}\times{}^{n_{x}}\to\real$ be a positive-definite kernel function. Consider the pre-Hilbert space of functions
${\cal H}_{K}^{0}({}^{n_{x}})=\left\{v:{}^{n_{x}}\to\real~{}\bigg{|}~{}\exists\,m\in\mathbb{N},\,{\boldsymbol{\omega}}\in{}^{m},\,\left\{\mathbf{x}\right\}_{j=1}^{m}\subset{}^{n_{x}}~{}\textup{such that}~{}v(\mathbf{x})=\sum_{j=1}^{m}\omega_{j}K(\mathbf{x}_{j},\mathbf{x})\right\}.$
The reproducing kernel Hilbert space (RKHS) ${\cal H}_{K}({}^{n_{x}})$ induced by the kernel $K$ is the (unique) completion of ${\cal H}_{K}^{0}({}^{n_{x}})$ with respect to the norm $\left\|\cdot\right\|_{{\cal H}_{K}({}^{n_{x}})}\coloneqq\left\langle\cdot,\cdot\right\rangle_{{\cal H}_{K}({}^{n_{x}})}^{1/2}$ induced by the inner product
$\left\langle v,v^{\prime}\right\rangle_{{\cal H}_{K}({}^{d})}\coloneqq\sum_{j=1}^{m}\sum_{k=1}^{m^{\prime}}\omega_{j}\omega_{k}^{\prime}K(\mathbf{x}_{j},\mathbf{x}_{k}^{\prime}),$
@@
For an ordered collection of pairwise-distinct vectors $\{\mathbf{x}_{j}\}_{j=1}^{m}\subset{}^{n_{x}}$, we use $K(\mathbf{X},\mathbf{X})$ to denote $m\times m$ kernel matrix of Definition 2.1 and define the vector $K(\mathbf{X},\mathbf{x})=[~{}K(\mathbf{x}_{1},\mathbf{x})~{}~{}\cdots~{}~{}K(\mathbf{x}_{m},\mathbf{x})~{}]^{\mathsf{T}}\in{}^{m}$. To simplify notation, we will write ${\cal H}_{K}$ for ${\cal H}_{K}({}^{n_{x}})$ when it is understood that $K$ is defined over ×${}^{n_{x}}\times{}^{n_{x}}$. Importantly, for $v(\mathbf{x})=\sum_{j=1}^{m}\omega_{j}K(\mathbf{x}_{j},\mathbf{x})$, the induced RKHS norm $\left\|v\right\|_{{\cal H}_{K}}$ can be computed efficiently via the corresponding kernel matrix,
@@
Let $v:{}^{n_{x}}\to\real$, $\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}\subset{}^{n_{x}}$ be pairwise distinct, and denote $y_{j}=v(\mathbf{x}_{j})$. For a given RKHS ${\cal H}_{K}$ and a regularization parameter $\gamma\geq 0$, a regularized interpolant $s_{v}^{\gamma}\in{\cal H}_{K}$ of $v$ is a solution to the minimization problem
@@
The minimization problem eq. 2.2 has a solution of the form
@@
where the coefficient vector ${\boldsymbol{\omega}}=[~{}\omega_{1}~{}~{}\cdots~{}~{}\omega_{m}~{}]^{\mathsf{T}}\in{}^{m}$ solves the $m\times m$ linear system
@@
Moreover, if $K$ is strictly positive definite, then $s_{v}^{\gamma}$ is the unique minimizer of eq. 2.2.
See, e.g., [50, Theorem 9.3] for a proof of Theorem 2.1. A key observation from Theorem 2.1 is that a solution to the infinite-dimensional minimization problem eq. 2.2 can be obtained by solving the finite-dimensional linear system eq. 2.3e.
Without regularization ($\gamma=0$), the function $s_{v}^{0}\in{\cal H}_{K}$ exactly interpolates the data, i.e., $s_{v}(\mathbf{x}_{j})=y_{j}$ for each $j=1,\dots,m$. Moreover, $s_{v}^{0}$ satisfies the following error bound.
@@
If $v\in{\cal H}_{K}$ and $s_{v}^{0}\in{\cal H}_{K}$ is an (unregularized) interpolant of $v$ corresponding to the pairwise distinct data $\left\{\mathbf{x}_{i}\right\}_{j=1}^{m}\subset{}^{n_{x}}$ and $y_{j}=v(\mathbf{x}_{j})\in\real$, then
$|v(\mathbf{x})-s_{v}^{0}(\mathbf{x})|\leq P_{K,\mathbf{X}}(\mathbf{x})\left\|v\right\|_{{\cal H}_{K}}\qquad\forall\;\mathbf{x}\in{}^{d},$ (2.4a)
where $P_{K,\mathbf{X}}:{}^{n_{x}}\to\real$ is the so-called power function defined by
@@
See, e.g., [49, Thm 4.9] for a proof of the bound eq. 2.4a and [49, Prop. 4.11, Prop. 4.12] for the characterization eq. 2.4b of the power function. While this error bound is for the unregularized, fully interpolatory case, it is still useful in practice for the regularized case when $\gamma>0$ is small.
@@ 2.2 Vector-valued kernel interpolation
Consider the vector-valued function $\mathbf{v}:{}^{n_{x}}\to{}^{n_{y}}$, $n_{y}>1$. As before, let $\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}\subset{}^{n_{x}}$ be pairwise distinct and suppose $\mathbf{y}_{j}=\mathbf{v}(\mathbf{x}_{j})\in{}^{n_{y}}$ for $j=1,\ldots,m$. Also let $v_{i}:{}^{n_{x}}\to\real$ denote the $i$-th component of $\mathbf{v}$, $y_{j,i}$ be the $i$-th component of $\mathbf{y}_{j}$, and define the input and output data matrices
$\mathbf{X}$ $=[~{}\mathbf{x}_{1}~{}~{}\cdots~{}~{}\mathbf{x}_{m}~{}]\in{}^{n_{x}\times m},$ $\mathbf{Y}$ $=[~{}\mathbf{y}_{1}~{}~{}\cdots~{}~{}\mathbf{y}_{m}~{}]\in{}^{n_{y}\times m}.$ (2.5)
@@
where each coefficient vector ${\boldsymbol{\omega}}_{1},\ldots,{\boldsymbol{\omega}}_{n_{y}}\in{}^{m}$ solves an $m\times m$ linear system,
@@
The minimization problem eq. 2.6 has a solution of the form
@@
where the coefficient matrix ${\boldsymbol{\Omega}}\in{}^{m\times{n_{y}}}$ solves the linear system
@@
Let $\mathbf{v}\in{\cal H}_{K}^{n_{y}}$ and $\mathbf{M}\in{}^{{n_{y}}\times{n_{y}}}$ be a symmetric positive definite weighting matrix with Cholesky factorization $\mathbf{M}=\mathbf{L}\mathbf{L}^{\mathsf{T}}$. If $\mathbf{s}_{\mathbf{v}}^{0}\in{\cal H}_{K}^{n_{y}}$ is an (unregularized) vector-valued interpolant of $\mathbf{v}$ of the form eq. 2.9 corresponding to the pairwise distinct data $\left\{\mathbf{x}_{i}\right\}_{i=1}^{m}\subset{}^{n_{x}}$ and $\mathbf{y}_{i}=\mathbf{v}(\mathbf{x}_{i})\in{}^{n_{y}}$, then
$\left\|\mathbf{v}(\mathbf{x})-\mathbf{s}_{\mathbf{v}}^{0}(\mathbf{x})\right\|_{\mathbf{M}}\leq P_{K,\mathbf{X}}(\mathbf{x})\left\|\mathbf{L}\right\|_{2}\left\|\mathbf{v}\right\|_{{\cal H}_{K}^{n_{y}}}\qquad\forall\;\mathbf{x}\in{}^{n_{x}}.$ (2.10)
@@ 2.3 Kernel selection
Since a positive-definite kernel $K$ uniquely defines the RKHS ${\cal H}_{K}$, the choice of kernel determines what form an interpolant can take as well as the approximation power of the optimal interpolant. We argue for the use of different types of kernels depending on how much information is available about the function $\mathbf{v}:{}^{n_{x}}\to{}^{n_{y}}$ being interpolated.
@@ 2.3.1 Unknown structure: radial basis function kernels
where $\psi:{}_{\geq 0}\to\real$ and $\epsilon>0$. Hence, RBF kernel interpolants are given by
$\mathbf{s}_{\mathbf{v}}^{\gamma}(\mathbf{x})={\boldsymbol{\Omega}}^{\mathsf{T}}{\boldsymbol{\psi}}_{\!\epsilon}(\mathbf{x}),\qquad{\boldsymbol{\psi}}_{\!\epsilon}(\mathbf{x})=\begin{bmatrix}\psi(\epsilon\left\|\mathbf{x}_{1}-\mathbf{x}\right\|_{2})\\ \vdots\\ \psi(\epsilon\left\|\mathbf{x}_{m}-\mathbf{x}\right\|_{2})\end{bmatrix}\in{}^{m}.$ (2.11b)

### v2

@@ 1 Introduction
Our proposed kernel-based non-intrusive ROMs, which we call “Kernel ROMs”, share similarities with the aforementioned approaches while overcoming some noticeable drawbacks. Like other approaches, we begin by applying POD or QM dimension reduction to a set of training snapshots and learn a function that approximates the system dynamics within a latent space. However, instead of modeling the ROM dynamics as a polynomial and learning the polynomial coefficients through least-squares regression as in OpInf and LaSDI, we use regularized kernel interpolation [37, 50, 51] to model the reduced dynamics with a function belonging to a user-defined reproducing kernel Hilbert space (RKHS). The structure of the learned function depends on the positive-definite kernel that defines the RKHS. For example, if the governing FOM has a polynomial structure, we can use a kernel induced by a feature map to compute ROM dynamics that share the same polynomial structure. On the other hand, if the FOM dynamics have unknown or only partially known structure, a more generic nonlinear kernel can be used to model the unknown part of the ROM dynamics. In this sense, our proposed approach has a natural way of incorporating closure terms into the ROM dynamics. While kernel methods have been used to emulate reduced dynamics in previous work [42, 63], these approaches are intrusive in that they assume that the FOM dynamics can be sampled explicitly, and they do not demonstrate a way to inject explicit structure into the learned ROM. The authors in [3] propose a generalization of DMD using kernel methods to learn a high-dimensional dynamical system consisting of a linear term and a nonlinear term, implicitly injecting structure by modeling nonlinear terms using polynomial kernels. The structure is implicit because a post-processing step is required to extract the linear term. While this approach bears some similarities to ours, we focus on constructing low-dimensional ROMs by applying a dimension reduction technique, projecting the training snapshots onto a low-dimensional latent space, and learning a low-dimensional dynamical system in the latent space whose structure explicitly mirrors the FOM structure. The key to preserving the FOM structure is the use of feature map kernels, which enables identifying the ROM coefficient matrices without requiring a post-processing step. In summary, our proposed approach is entirely data-driven, can produce interpretable and flexible ROMs, and yields computable a posteriori error bounds between the non-intrusive ROM and FOM solutions.
@@ 2.1 Scalar-valued kernel interpolation
We begin with a review of regularized kernel interpolation for scalar-valued functions. By the Moore–Aronszajn Theorem (see, e.g., [50, Theorem 3.10]), a positive-definite kernel function defines a unique Hilbert space with desirable properties. This result leads to the Representer Theorem — the key result used for computing an optimal interpolant in an RKHS — as well as a pointwise error bound on the interpolant.
@@
A function $K:\real^{n_{x}}\times\real^{n_{x}}\to\real$ is a (real-valued) kernel function if it is symmetric, i.e., $K(\mathbf{x},\mathbf{x}^{\prime})=K(\mathbf{x}^{\prime},\mathbf{x})$ for all $\mathbf{x},\mathbf{x}^{\prime}\in\real^{n_{x}}$. A kernel function $K$ is said to be positive definite if for any matrix $\mathbf{X}=[~\mathbf{x}_{1}~~\cdots~~\mathbf{x}_{m}~]\in\real^{n_{x}\times m}$ with pairwise distinct columns, the kernel matrix $K(\mathbf{X},\mathbf{X})\in\real^{m\times m}$ with entries $K(\mathbf{X},\mathbf{X})_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$ is positive semi-definite. If $K(\mathbf{X},\mathbf{X})$ is strictly positive definite, then $K$ is said to be strictly positive definite.
@@
Let $K:\real^{n_{x}}\times\real^{n_{x}}\to\real$ be a positive-definite kernel function. Consider the pre-Hilbert space of functions
${\cal H}_{K}^{0}(\real^{n_{x}})=\left\{v:\real^{n_{x}}\to\real~\bigg|~\exists\,m\in\mathbb{N},\,{\boldsymbol{\omega}}\in\real^{m},\,\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}\subset\real^{n_{x}}~\textup{such that}~v(\mathbf{x})=\sum_{j=1}^{m}\omega_{j}K(\mathbf{x}_{j},\mathbf{x})\right\}.$
The reproducing kernel Hilbert space (RKHS) ${\cal H}_{K}(\real^{n_{x}})$ induced by the kernel $K$ is the (unique) completion of ${\cal H}_{K}^{0}(\real^{n_{x}})$ with respect to the norm $\left\|\cdot\right\|_{{\cal H}_{K}(\real^{n_{x}})}\coloneqq\left\langle\cdot,\cdot\right\rangle_{{\cal H}_{K}(\real^{n_{x}})}^{1/2}$ induced by the inner product
$\left\langle v,v^{\prime}\right\rangle_{{\cal H}_{K}(\real^{d})}\coloneqq\sum_{j=1}^{m}\sum_{k=1}^{m^{\prime}}\omega_{j}\omega_{k}^{\prime}K(\mathbf{x}_{j},\mathbf{x}_{k}^{\prime}),$
@@
For an ordered collection of pairwise-distinct vectors $\{\mathbf{x}_{j}\}_{j=1}^{m}\subset\real^{n_{x}}$, we use $K(\mathbf{X},\mathbf{X})$ to denote $m\times m$ kernel matrix of Definition 2.1 and define the vector $K(\mathbf{X},\mathbf{x})=[~K(\mathbf{x}_{1},\mathbf{x})~~\cdots~~K(\mathbf{x}_{m},\mathbf{x})~]^{\mathsf{T}}\in\real^{m}$. To simplify notation, we will write ${\cal H}_{K}$ for ${\cal H}_{K}(\real^{n_{x}})$ when it is understood that $K$ is defined over $\real{}^{n_{x}}\times\real^{n_{x}}$. Importantly, for $v(\mathbf{x})=\sum_{j=1}^{m}\omega_{j}K(\mathbf{x}_{j},\mathbf{x})$, the induced RKHS norm $\left\|v\right\|_{{\cal H}_{K}}$ can be computed efficiently via the corresponding kernel matrix,
@@
Let $v:\real^{n_{x}}\to\real$, $\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}\subset\real^{n_{x}}$ be pairwise distinct, and denote $y_{j}=v(\mathbf{x}_{j})$. For a given RKHS ${\cal H}_{K}$ and a regularization parameter $\gamma\geq 0$, a regularized interpolant $s_{v}^{\gamma}\in{\cal H}_{K}$ of $v$ is a solution to the minimization problem
@@
If $\gamma>0$, then the minimization problem eq. 2.2 has a solution of the form
@@
where the coefficient vector ${\boldsymbol{\omega}}=[~\omega_{1}~~\cdots~~\omega_{m}~]^{\mathsf{T}}\in\real^{m}$ solves the $m\times m$ linear system
@@
Moreover, if $K$ is strictly positive definite, then $s_{v}^{\gamma}$ is the unique minimizer of eq. $2.$
See, e.g., [51, Theorem 9.3] for a proof of Theorem 2.1. A key observation from Theorem 2.1 is that a solution to the infinite-dimensional minimization problem eq. 2.2 can be obtained by solving the finite-dimensional linear system eq. 2.3e.
Without regularization ($\gamma=0$) and assuming that $K(\mathbf{X},\mathbf{X})$ is invertible, the function $s_{v}^{0}\in{\cal H}_{K}$ exactly interpolates the data, i.e., $s_{v}(\mathbf{x}_{j})=y_{j}$ for each $j=1,\dots,m$. Moreover, $s_{v}^{0}$ satisfies the following error bound.
@@
If $v\in{\cal H}_{K}$ and $s_{v}^{0}\in{\cal H}_{K}$ is an (unregularized) interpolant of $v$ corresponding to the pairwise distinct data $\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}\subset\real^{n_{x}}$ and $y_{j}=v(\mathbf{x}_{j})\in\real$, then
$|v(\mathbf{x})-s_{v}^{0}(\mathbf{x})|\leq P_{K,\mathbf{X}}(\mathbf{x})\left\|v\right\|_{{\cal H}_{K}}\qquad\forall\;\mathbf{x}\in\real^{d},$ (2.4a)
where $P_{K,\mathbf{X}}:\real^{n_{x}}\to\real$ is the so-called power function defined by
@@
See, e.g., [50, Thm 4.9] for a proof of the bound eq. 2.4a and [50, Prop. 4.11, Prop. 4.12] for the characterization eq. 2.4b of the power function. While this error bound is for the unregularized, fully interpolatory case, it is still useful in practice for the regularized case when $\gamma>0$ is small.
@@ 2.2 Vector-valued kernel interpolation
Consider the vector-valued function $\mathbf{v}:\real^{n_{x}}\to\real^{n_{y}}$, $n_{y}>1$. As before, let $\left\{\mathbf{x}_{j}\right\}_{j=1}^{m}\subset\real^{n_{x}}$ be pairwise distinct and suppose $\mathbf{y}_{j}=\mathbf{v}(\mathbf{x}_{j})\in\real^{n_{y}}$ for $j=1,\ldots,m$. Also let $v_{i}:\real^{n_{x}}\to\real$ denote the $i$-th component of $\mathbf{v}$, $y_{j,i}$ be the $i$-th component of $\mathbf{y}_{j}$, and define the input and output data matrices
$\mathbf{X}$ $=[~\mathbf{x}_{1}~~\cdots~~\mathbf{x}_{m}~]\in\real^{n_{x}\times m},$ $\mathbf{Y}$ $=[~\mathbf{y}_{1}~~\cdots~~\mathbf{y}_{m}~]\in\real^{n_{y}\times m}.$ (2.5)
@@
where each coefficient vector ${\boldsymbol{\omega}}_{1},\ldots,{\boldsymbol{\omega}}_{n_{y}}\in\real^{m}$ solves an $m\times m$ linear system,
@@
If $\gamma>0$, then the minimization problem eq. 2.6 has a solution of the form
@@
where the coefficient matrix ${\boldsymbol{\Omega}}\in\real^{m\times{n_{y}}}$ solves the linear system
@@
Let $\mathbf{v}\in{\cal H}_{K}^{n_{y}}$ and $\mathbf{M}\in\real^{{n_{y}}\times{n_{y}}}$ be a symmetric positive definite weighting matrix with Cholesky factorization $\mathbf{M}=\mathbf{L}\mathbf{L}^{\mathsf{T}}$. If $\mathbf{s}_{\mathbf{v}}^{0}\in{\cal H}_{K}^{n_{y}}$ is an (unregularized) vector-valued interpolant of $\mathbf{v}$ of the form eq. 2.9 corresponding to the pairwise distinct data $\left\{\mathbf{x}_{i}\right\}_{i=1}^{m}\subset\real^{n_{x}}$ and $\mathbf{y}_{i}=\mathbf{v}(\mathbf{x}_{i})\in\real^{n_{y}}$, then
$\left\|\mathbf{v}(\mathbf{x})-\mathbf{s}_{\mathbf{v}}^{0}(\mathbf{x})\right\|_{\mathbf{M}}\leq P_{K,\mathbf{X}}(\mathbf{x})\left\|\mathbf{L}\right\|_{2}\left\|\mathbf{v}\right\|_{{\cal H}_{K}^{n_{y}}}\qquad\forall\;\mathbf{x}\in\real^{n_{x}}.$ (2.10)
@@ 2.3 Kernel selection
Since a positive-definite kernel $K$ uniquely defines the RKHS ${\cal H}_{K}$, the choice of kernel determines what form an interpolant can take as well as the approximation power of the optimal interpolant. We argue for the use of different types of kernels depending on how much information is available about the function $\mathbf{v}:\real^{n_{x}}\to\real^{n_{y}}$ being interpolated.
@@ 2.3.1 Unknown structure: radial basis function kernels
where $\psi:\real_{\geq 0}\to\real$ and $\epsilon>0$. Hence, RBF kernel interpolants are given by
$\mathbf{s}_{\mathbf{v}}^{\gamma}(\mathbf{x})={\boldsymbol{\Omega}}^{\mathsf{T}}{\boldsymbol{\psi}}_{\!\epsilon}(\mathbf{x}),\qquad{\boldsymbol{\psi}}_{\!\epsilon}(\mathbf{x})=\begin{bmatrix}\psi(\epsilon\left\|\mathbf{x}_{1}-\mathbf{x}\right\|_{2})\\ \vdots\\ \psi(\epsilon\left\|\mathbf{x}_{m}-\mathbf{x}\right\|_{2})\end{bmatrix}\in\real^{m}.$ (2.11b)